OpenStudy (anonymous):

What is the equation of the line that passes through the points (-3, 2) and (-5, 8)?

OpenStudy (anonymous):

Slope = 8-2 / -5 - 3 = 6/-8 = -3/4 y - y1 = m (x-x1) y - 2 = -3/4(x + 3) y - 2 = -3/4x - 9/4 y = -3/4x -9/4 + 2 y = -3/4x -9/4 + 8/4 y = -3/4x - 1/4